a cyclist traveled 60 mi at a constant rate before reducing the speed by 7mph. Another 40 mi was traveled at the reduced speed. The total time for the 100-mile trip was 9 h. Find the rate during the first 60 mi.
Answer by stanbon(75887)

a cyclist traveled 60 mi at a constant rate before reducing the speed by 7mph. Another 40 mi was traveled at the reduced speed. The total time for the 100-mile trip was 9 h. Find the rate during the first 60 mi.
-----
Original leg DATA:
dist = 60 miles ; rate = x mph; time = 60/x hrs
----------------------
2nd leg DATA:
distance = 40 miles ; rate = x-7 mph ; time = 40/(x-7) hrs
-------------------------
Equation:
time + time = 9 hrs
-----
60/x + 40/(x-7) = 9
60(x-7) + 40x = 9x(x-7)
------
100x - 420 = 9x^2 - 63x
-----
9x^2 - 163x + 420 = 0
x = 15 mph
======================
Cheers,
Stan H.
-----------------------